Please answer my questions.


1) Calculate the total no. of electrons present in one mole of methane.


2) FInd (a) the total no. and (b) the total mass of neutrons in 7mg of  14 c.


  [Assume mass of the neutron  = 1.675 x 10-2 7kg]


3)Find (a) the total no. and (b) the total mass of protons in 34 mg of NH3 at STP.


    Will the answer change if the temperature and pressure are changed?


Please answer as quick as possible..

Grade:12th pass

4 Answers

Gurwinder Kaur
12 years ago
(i)number of electrons present in 1 mole of methane(CH4); {1(6)+4(1)}=10 Number of electrons present in 1 mole i.e.6.023*10 power 23 molecules of methane=6.022*10 power 23*10=6.022*10 power 24 (ii) (a)number of atoms of 14 c in one mole 6.023*10 power 23 Since 1 atom of 14 c contains (14-6) i.e., 8 neutrons,the no. of neutrons in 14g of 14 c is (6.023*10 power 23)*8. Or,14g of 14 c contains (6.022*10 power 23*8)neutrons. Numbers of neutrons in 7mg = (6.022*10 power 23*8*7mg)/1400mg= 2.4092*10 power 21 (b) Mass of 1 neutron = 1.67493*10 power -27kg Mass of total neutrons in 7g of 14 c = (2.4092*10 power 21)(1.67493*10 power -27)=4.0352*10 power -6 kg (iii) (a) 1 mole of NH3={1(14)+3(1)}=17g if NH3=6.022*10 power 23 molecules of NH3 Total no. of Protons present in 1 mole of NH3={1(7)+3(1)}=10 no. of protons in 6.022*10 power 23 molecules of NH3=(6.022*10 power 23)*10=6.022*10 power 24 17g of NH3 contains (6.022*10 power 24)protons Number of protons in 34mg of NH3=(6.022*10 power 24*34mg)/17000mg=1.2046*10 power 22 (b) mass of 1 proton =1.67493*10 power -27kg total mass of proton in 34mg of NH3=(1.67493*10 power -27kg)(1.2046*10 power 22)=2.0176*10 power -5kg The no. of protons, neutrons & electrons in an atom is independent of temperature and pressure conditions.Hence the obtained value will remained unchanged if the temprature and pressure is changed.
Gurwinder Kaur
12 years ago

(i)number of electrons present in 1 mole of methane(CH4); {1(6)+4(1)}=10 Number of electrons present in 1 mole i.e.6.023*10 power 23 molecules of methane=6.022*10 power 23*10=6.022*10 power 24 (ii) (a)number of atoms of 14 c in one mole 6.023*10 power 23 Since 1 atom of 14 c contains (14-6) i.e., 8 neutrons,the no. of neutrons in 14g of 14 c is (6.023*10 power 23)*8. Or,14g of 14 c contains (6.022*10 power 23*8)neutrons. Numbers of neutrons in 7mg = (6.022*10 power 23*8*7mg)/1400mg= 2.4092*10 power 21 (b) Mass of 1 neutron = 1.67493*10 power -27kg Mass of total neutrons in 7g of 14 c = (2.4092*10 power 21)(1.67493*10 power -27)=4.0352*10 power -6 kg (iii) (a) 1 mole of NH3={1(14)+3(1)}=17g if NH3=6.022*10 power 23 molecules of NH3 Total no. of Protons present in 1 mole of NH3={1(7)+3(1)}=10 no. of protons in 6.022*10 power 23 molecules of NH3=(6.022*10 power 23)*10=6.022*10 power 24 17g of NH3 contains (6.022*10 power 24)protons Number of protons in 34mg of NH3=(6.022*10 power 24*34mg)/17000mg=1.2046*10 power 22 (b) mass of 1 proton =1.67493*10 power -27kg total mass of proton in 34mg of NH3=(1.67493*10 power -27kg)(1.2046*10 power 22)=2.0176*10 power -5kg The no. of protons, neutrons & electrons in an atom is independent of temperature and pressure conditions.Hence the obtained value will remained unchanged if the temprature and pressure is changed.

Gurwinder Kaur
12 years ago

(i)number of electrons present in 1 mole of methane(CH4); {1(6)+4(1)}=10

 Number of electrons present in 1 mole i.e.6.023*10 power 23 molecules of methane=6.022*10 power 23*10=6.022*10 power 24

 (ii) (a)number of atoms of 14 c in one mole 6.023*10 power 23

Since 1 atom of 14 c contains (14-6) i.e., 8 neutrons,the no. of neutrons in 14g of 14 c is (6.023*10 power 23)*8. Or,14g of 14 c contains (6.022*10 power 23*8)neutrons.

 Numbers of neutrons in 7mg = (6.022*10 power 23*8*7mg)/1400mg= 2.4092*10 power 21

(b) Mass of 1 neutron = 1.67493*10 power -27kg Mass of total neutrons in 7g of 14 c = (2.4092*10 power 21)(1.67493*10 power -27)=4.0352*10 power -6 kg

(iii) (a) 1 mole of NH3={1(14)+3(1)}=17g of NH3=6.022*10 power 23 molecules of NH3

Total no. of Protons present in 1 mole of NH3={1(7)+3(1)}=10

 no. of protons in 6.022*10 power 23 molecules of NH3=(6.022*10 power 23)*10=6.022*10 power 24

17g of NH3 contains (6.022*10 power 24)protons

 Number of protons in 34mg of NH3=(6.022*10 power 24*34mg)/17000mg=1.2046*10 power 22

 (b) mass of 1 proton =1.67493*10 power -27kg

total mass of proton in 34mg of NH3=(1.67493*10 power -27kg)(1.2046*10 power 22)=2.0176*10 power -5kg

The no. of protons, neutrons & electrons in an atom is independent of temperature and pressure conditions.Hence the obtained value will remained unchanged if the temprature and pressure is changed.

plz approve by clicking yes below if it is helpful....

Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to the problem below.
 
(1) number of electrons present in 1 mole of methane(CH4); {1(6)+4(1)} = 10
Number of electrons present in 1 mole i.e.6.023*1023 molecules of methane = 6.022 x 1023 x 10 = 6.022 x 1024
 
(2) (a) number of atoms of  C – 14 in one mole 6.023*1023
Since 1 atom of 14 c contains (14 – 6) i.e., 8 neutrons, the no. of neutrons in 14g of C – 14 is (6.023*1023)*8. Or,14g
of C – 14 contains (6.022*1023 * 8)neutrons.
Numbers of neutrons in 7mg = (6.022*1023 x 8 x 7mg)/1400mg = 2.4092 x 1021
 
(b) Mass of 1 neutron = 1.67493*10-27kg Mass of total neutrons in 7g of 14 c = (2.4092*1021)(1.67493*10-27)=4.0352*10-6 kg
 
(3) (a) 1 mole of NH3 = {1(14)+3(1)} = 17g of NH3 = 6.022*1023 molecules of NH3
Total no. of Protons present in 1 mole of NH3 = {1(7)+3(1)} = 10
no. of protons in 6.022*10 power 23 molecules of NH3 = (6.022*1023)*10 = 6.022*1024
17g of NH3 contains (6.022*1024)protons
Number of protons in 34mg of NH3 = (6.022*1024*34mg)/17000mg=1.2046*1022
 
(b) mass of 1 proton =1.67493*10-27kg
total mass of proton in 34mg of NH3=(1.67493*10-27kg)(1.2046*1022)=2.0176*10-5kg
The no. of protons, neutrons & electrons in an atom is independent of temperature and pressure conditions.Hence the
obtained value will remained unchanged if the temperature and pressure is changed.
 
Hope it helps.
Thanks and regards,
Kushagra